OpenStudy (anonymous):

I need help with this problem. If a * b = a2 -b 2. Find 3 * 2. ________ a + b

OpenStudy (anonymous):

I' m attaching a file with 5 choices for this problem.

jimthompson5910 (jim_thompson5910):

Hint: \[\Large a * b = \frac{a^2-b^2}{a+b}\] \[\Large 3 * 2 = \frac{3^2-2^2}{3+2}\]

OpenStudy (anonymous):

so it would be -2

jimthompson5910 (jim_thompson5910):

not quite

OpenStudy (anonymous):

2 then

jimthompson5910 (jim_thompson5910):

\[\Large 3 * 2 = \frac{3^2-2^2}{3+2}\] \[\Large 3 * 2 = \frac{9-4}{3+2}\] \[\Large 3 * 2 = \frac{5}{5}\] \[\Large 3 * 2 = 1\]